OpenStudy (anonymous):

3(2+3)-7+8/2 I got 4, but the book is telling me its 12. help please?

OpenStudy (gw2011):

(3)(5)-7+4=15-7+4=12

OpenStudy (anonymous):

6+9-7+4 15-7+4 12

OpenStudy (anonymous):

Assuming Muse knows basic fact, the question is the order of evaluation. Division before addition. Everything within parentheses before anything outside. Muse, how did you compute 4?

OpenStudy (anonymous):

heres what I did: 3(2+3)-7+8/2 3(5)-7+8/2 15-7+8/2 15-7+4 15-11=4

OpenStudy (anonymous):

distribute 3 to the 2 and 3 within the parenthesis. you will get 6+9-7+8/2. afterwards, divide 8 by 2. you will get 4 as an answer; 6+9-7+4. solve left to right and you will get: 6+9=15 15-7=8 8+4=12 12 is the correct answer.

OpenStudy (anonymous):

oh I see. so I added 7+4 when I was suppose to actually subtract 15-11 first.

OpenStudy (anonymous):

was I suppose to follow the PEMDAS rule on this problem, or did it not need apply?